OpenStudy (anonymous):

I am trying to help my son with his work...the question is...27x(1/3)3...the 3 is to the 3rd power

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

\[27*(\frac{1}{3})^3=27*\frac{1}{3^3}=27*\frac{1}{27}=\frac{27}{27}=1\]

OpenStudy (angela210793):

(1/3)^3=1/27 27*1/27=1

OpenStudy (anonymous):

thank you